OpenStudy (anonymous):

integral x^3/(x^3 +1)

OpenStudy (jamesj):

This is a bit ugly. Remember that x^3 + 1 = (x+1)(x^2 - x + 1) This points you down the partial fractions route, and that's where you need to go.

OpenStudy (anonymous):

i think you can do this by the substitution u = x^3

OpenStudy (anonymous):

so far i've gotten it to: x -1/3ln{x+1} + integral of (-x/3 + 2/3)/(x^2 -x +1), is that right? and if so how do i do i integrate it from there

OpenStudy (anonymous):

You can first simplify it to \(1-\frac{1}{x^3+1}=1-\frac{1}{(x+1)(x^2-x+1)}\). Then you should use partial fractions to expand the second term.

OpenStudy (jamesj):

So x^3/(x^3 + 1) = 1 -1/(x^3+1) so the integrating the RHS, there is definitely an x. Now we need to write down the partial fraction expansion of the 1/(x^3+1), which is \[\frac{2-x}{3(x^2 - x +1)} + \frac{1}{3(x+1)}\] So yes, I agree with you so far. Now you separate the two terms over x^2 - x + 1. One will give you a log, the other an arctan.

OpenStudy (jamesj):

just be careful with the signs: I might have made an error, or you've made an error as there's a small inconsistency between us.

OpenStudy (anonymous):

how would i do the integral of (-x/3 + 2/3)/(x^2 -x +1

OpenStudy (jamesj):

the integral of x/(x^2 - x + 1) = x/((x-1/2)^2 + (sqrt(3)/2)^2) Now you make a substitution of u = x - 1/2 The integral of 1/((x^2 - x + 1) = 1/((x-1/2)^2 + (sqrt(3)/2)^2) is an arctan: set x = 1/2 + sqrt(3)/2.tan u

OpenStudy (anonymous):

but what happened to the -1/3 and 2/3?

OpenStudy (jamesj):

I took them out just to simply the description, but obviously you'll need to put them back in an scale back up the integrals.

OpenStudy (jamesj):

...just to simplify...

OpenStudy (anonymous):

could i just multiply the equation by 3 to get rid of the rations?

OpenStudy (jamesj):

Sure ... whatever makes it easier for you. Just remember to scale it back at the end.

OpenStudy (anonymous):

i don't see how you broke up the denominator though?

OpenStudy (jamesj):

\[x^2 - x + 1 = x^2 - x + 1/4 + 3/4 = (x - 1/2)^2 + (\sqrt{3}/2)^2\]